# Fluorescence emission of the JUNO liquid scintillator

### v2

@@ 2 Composition and properties of linear alkylbenzene based liquid scintillators
The composition of the scintillator mixture determines its optical properties, including the time distribution of fluorescence light, the light yield and the emission and absorption spectra.
@@ 3.2 Measurements
We measured the emission spectrum of the JUNO liquid scintillator between 250 nm and 600 nm in a 1 cm-path length cuvette in a right angle geometry, as shown in figure 1 (figure 2 shows a zoom between 300 nm and 550 nm). One can see how the emission spectrum changes when a component is added to the LAB. The violet, green and black lines show the emission spectra of LAB + 2.5 g/L PPO + 3 mg/L Bis-MSB (the recipe of the JUNO liquid scintillator), LAB + 2.5 g/L PPO and pure LAB. The final spectrum reported here is considered only in a small geometry and it could change by the one seen by the JUNO PMTs due to the much longer propagation.
@@ 4 Absorbance and absorption length
Note that given the reduced length of our setup cuvettes, our sensitivity on the absorption lengths is limited to $\leq$ 1 m. However, considering the rapid increasing trend at $\sim$ 405 nm, we expect the absorption length to be $\geq$ 18 m for wavelengths $\geq$ 430 nm, consistent with the results observed in liquid scintillators of similar composition and similar solvent [23, 24, 25].
@@ 5.3 Data analysis
In order to evaluate the validity of the four exponential model we also tried to fit our data assuming a model based on three and five exponentials. In the former case we observed a worse $\chi^{2}$, while in the latter one the fifth component contribution is zero. This leads us to the conclusion that the four exponential model is the best option to describe our data.

### v4

@@ 2 Composition and properties of linear alkylbenzene based liquid scintillators
The amount of flours in the scintillator mixture determines its optical properties, including the time distribution of fluorescence light, the light yield and the emission and absorption spectra.
@@ 3.2 Measurements
We measured the emission spectrum of the JUNO liquid scintillator between 250 nm and 600 nm in a 1 cm-path length cuvette in a right angle geometry, as shown in figure 1 (figure 2 shows a zoom between 300 nm and 550 nm). One can see how the emission spectrum changes when a component is added to the LAB. The violet, green and black lines show the emission spectra of LAB + 2.5 g/L PPO + 3 mg/L Bis-MSB (the recipe of the JUNO liquid scintillator), LAB + 2.5 g/L PPO and pure LAB. We note that the spectrum reported here has been obtained in a small geometry setup ($\sim$ centimenters). In JUNO, due to its large dimensions ($\sim$ several tens of meters) the situation will be significantly different: in particular, the lower part of the spectrum will be absorbed and re-emitted above 400 nm. This would make the peaks at 400 and 420 more dominant and the total emission more similar to the bis-MSB one [19].
@@ 4 Absorbance and absorption length
Note that given the reduced length of our setup cuvettes, our sensitivity on the absorption lengths is limited to $\leq$ 1 m. However, considering the rapid increasing trend at $\sim$ 405 nm, we expect the absorption length to be $\geq$ 18 m for wavelengths $\geq$ 430 nm, consistent with the results observed in liquid scintillators of similar composition and similar solvent [23, 24, 25]. In addition, due to aging it is possible that the optical properties, in particular the attenuation length, will change over time. For this reason some experiments have been conducted, showing that in 4 years the LS performances are preserved with small effects on the attenuation length [26]. Therefore for a long living experiment like JUNO a careful calibration of the evolution of the optical properties of the scintillator inside the detector is mandatory.
@@ 5.3 Data analysis
In order to evaluate the validity of the four exponential model we also tried to fit our data assuming a model based on three and five exponentials. In the former case we observed a worse $\chi^{2}$, while in the latter one the fifth component contribution railed to zero. Furthermore,it is important to notice that the IRF parameters were fixed during the fitting procedure of the fluorescence curve to the value obtained by analyzing the IRF curve.